# An Efficient and Adaptive Next Edit Suggestion Framework with Zero Human Instructions in IDEs

### v1

@@ Abstract.
Code editing, including modifying, refactoring, and maintaining existing code, is the most frequent task in software development and has garnered significant attention from AI-powered tools. However, existing solutions that translate explicit natural language instructions into code edits face critical limitations, such as heavy reliance on human instruction input and high latency, which hinder their effective integration into a developer’s workflow.
We observe that developers’ habitual behaviors and coding objectives are often reflected in their historical editing patterns, making this data key to addressing existing limitations. To leverage these insights, we propose NES (Next Edit Suggestion), an LLM-driven code editing framework that delivers an instruction-free and low-latency experience. Built on a dual-model architecture and trained with our high-quality SFT and DAPO datasets, NES enhances productivity by understanding developer intent while optimizing inference to minimize latency.
NES is a scalable, industry-ready solution with a continuous Tab key interaction workflow, seamlessly adopted by a FinTech company with over 20,000 developers. Evaluations on real-world datasets show NES achieves 75.6% and 81.6% accuracy in two tasks of predicting next edit locations, alongside 91.36% ES and 27.7% EMR for intent-aligned edits, outperforming SOTA models. Our open-sourced SFT and DAPO datasets have been demonstrated to enhance the performance of open-source CodeLLMs. The demonstration of NES is available at https://youtu.be/yGoyYOe6fbY.
Code Edit, Reinforcement Learning, Large Language Models, Software Engineering
ccs: Software and its engineering
@@ 1. Introduction
With Large Language Models (LLMs) demonstrating remarkable capabilities in various aspects of software development (Di et al., 2023; Yang et al., 2025; Wang et al., 2024b; Guo et al., 2024; Liang et al., 2024; Rozière et al., 2023; Ding et al., 2024; Codefuse et al., 2025), AI-powered development tools such as GitHub Copilot (Copilot, 2023), Cursor (Cursor, 2024) and Zeta (Zed, 2025) are revolutionizing modern software development workflows by leveraging LLMs for code editing tasks (Wang et al., 2021; Zhang et al., 2022b; Gupta et al., 2023; Wei et al., 2024; Liu et al., 2024a). Empirical studies further highlight that code editing, which includes activities such as modifying (Li et al., 2024; Guo et al., 2025), refactoring, and maintaining existing code, is one of the most frequent and critical activities throughout the lifecycle of a software project (Nguyen et al., 2013). The importance of efficient and accurate code editing cannot be overstated, as it directly impacts the productivity and quality of software development.
Existing research efforts have made significant advancements, particularly in translating explicit natural language instructions into code edits. However, these approaches face critical limitations that hinder their effective integration into a developer’s workflow. On one hand, most existing methods rely heavily on human instruction input to maintain a coherent reference chain, disrupting programming flow and increasing cognitive load. Additionally, tools that fail to capture users’ habitual coding behavior often generate suggestions misaligned with their natural coding style. On the other hand, code editing tasks are inherently time-sensitive, with users expecting feedback within 1s (Card, 2018). Existing tools, however, rely on computationally intensive processes (e.g., search, natural language understanding), creating a tradeoff between low-latency and precise suggestions. This significant latency undermines real-time interaction and degrades user experience. Overcoming these limitations can undoubtedly enhance code editing in real-world development. However, it necessitates modifying the method of acquiring cues, which can pose significant challenges in accurately identifying the editing location and determining the content within the desired timeframe.
@@
We propose NES (Next Edit Suggestion), an LLM-driven code editing framework designed for an instruction-free and low-latency (feel like zero-latency) experience. Comprising a synergistic dual-model architecture trained on the proposed Supervised Fine-Tuning (SFT) and Dynamic sAmpling Policy Optimization (DAPO) datasets, NES adapts to user habits, integrates seamlessly into workflows, and enhances productivity by implicitly learning developer intent.
@@
  • NES-Edit Model analyzes the developer’s editing history, including refactoring patterns and API call sequences, to generate personalized and precise code modifications for the current editing region.
To improve user experience, NES leverages Prefix Caching (PC) (Zheng et al., 2023) and Speculative Decoding (SD) (Leviathan et al., 2022; Chen et al., 2023) techniques to achieve substantial inference speedups.
NES provides a scalable, industry-ready solution through its continuous Tab key interaction workflow, enabling seamless adoption in large-scale environments. Deployed in a FinTech company with 20,000+ developers, it handles thousands of daily code changes effortlessly. By allowing developers to perform complex tasks like refactoring or API call chains through simple "TabTabTab " actions, NES eliminates manual instructions, reduces cognitive load, and minimizes context-switching, delivering an efficient and immersive coding experience.
We evaluate NES using extensive real-world datasets from both open-source and industrial projects, highlighting its practical effectiveness. Experimental results show that the NES-Location Model achieves 75.7% accuracy in predicting the next edit location, while the NES-Edit Model delivers a 27.7% exact match rate (EMR) for generating intent-aligned code edits, outperforming several state-of-the-art (SOTA) models. Moreover, NES reduces inference time to under 450ms by optimizing model performance on Nvidia L20 GPUs. Finally, training on our proposed SFT and DAPO datasets results in notable improvements on Qwen2.5-Coder-7B and Seed-Coder-8B-Instrcut models.
@@
  • We propose NES  [^1] NES mentioned in the paper, unless specifically noted otherwise, use the Qwen3-4B+SFT+DAPO model. See Section 5 for details., a dual-model, LLM-driven code editing framework that leverages historical editing patterns and operates in an instruction-free, low-latency paradigm, eliminating the need for explicit natural language human input.
  • We build the high-quality SFT and DAPO datasets [^2] We have submitted the datasets to GitHub and will make them publicly available soon at https://anonymous.link.for.review,which offer significant improvements in performance. These datasets have been instrumental in achieving accuracy and similarity scores that are dozens of times higher on code editing tasks across a variety of models.
  • We have successfully integrated and deployed NES on Nvidia L20 GPUs within a FinTech organization. Utilizing a comprehensive set of inference optimizations, NES has achieved low-latency performance, with response times under 450ms, thereby enhancing the user experience.
@@ 2. Motivating Example
The process involves manual navigation, repetitive edits, and significant cognitive effort to track changes. Although LLM-driven tools can assist with locating and making edits, instruction-based tools, such as CodeT5 (Wang et al., 2021), CoditT5 (Zhang et al., 2022b), Coeditor (Wei et al., 2024), and CoEdPilot (Liu et al., 2024a), disrupt the developer’s workflow by requiring precise commands and waiting for responses, introducing additional cognitive overhead. To address these challenges, an instruction-free tool is essential, which is where NES comes in.
@@
A tedious, error-prone manual process is streamlined into an efficient "TabTabTab " workflow. Powered by high-quality datasets, NES predicts both the location (NES-Location) and content (NES-Edit) of edits without requiring explicit instructions. This reduces cognitive load, minimizes context switching, accelerates inference and delivers an immersive development experience.
@@ 3.1. An Overview of NES Framework
  • Next Location Suggestion: The NES-Location Model, leveraging real-time updated historical editing trajectories ①, predicts the developer’s potential next edit location. Subsequently, the framework presents this predictive suggestion to the developer, who then has the option to either rapidly navigate to the suggested position by pressing the Tab key or to disregard the suggestion ②.
  • Code Edit Suggestion: Thereafter, the NES-Edit Model, integrating historical editing patterns with the current editing window’s context, generates real-time code editing suggestions ③. Developers have the flexibility to either accept these suggestions by pressing the Tab key or continue with independent edits ④, while also updating the edit history ⑤.
  • Dynamic loop: Upon the developer’s adoption of an editing suggestion, the framework promptly initiates the next round of next location suggestion, establishing a continuous intelligent assistance loop. Ideally, developers can significantly enhance their coding efficiency and quality by consistently accepting NES framework’s suggestions by pressing the Tab key ⑥.
@@
  • Incremental Difference Detection: We developed a real-time code change monitor based on incremental computation. This framework accurately captures each developer editing operation, thereby constructing a comprehensive and detailed historical editing trajectory. These trajectory data serve as crucial inputs for subsequent model training and prediction processes.
  • High-Quality Training Dataset Construction: Utilizing the collected authentic historical editing trajectories, we constructed a high-quality training dataset. This dataset encompasses not only sequences of editing operations but also contextual information for each edit. Therefore, this dataset can enabling the training of both the NES-Location Model and the NES-Edit Model with nuanced understanding of editing patterns and their surrounding code environments.
  • Model Training: To endow NES with the capability for location and code editing suggestion, we designed a two-stage training methodology. Initially, both the NES-Location and NES-Edit models undergo SFT (Ouyang et al., 2022) on our large-scale historical editing datasets, enabling them to learn the fundamental patterns of code modification. Subsequently, we employ Decoupled Clip and DAPO (Yu et al., 2025), a reinforcement learning technique using high-quality preference data, to further refine the models. This second stage aligns their behavior more closely with real-world developer intent and utility, significantly boosting the practical value of NES.
  • Model Inference Optimization: Acknowledging the constraints of inference resources in practical application environments and the stringent low-latency requirements inherent in real-time editing suggestion, we conducted in-depth optimization of the model inference process, and achieved real-time responsiveness to developer editing behaviors, ensuring the fluidity and practicality of the NES framework.
@@ 3.2. Incremental Difference Detection
We developed an efficient incremental difference detection mechanism built upon our AI-powered IDE, designed to accurately capture real-world code editing trajectories of developers during real-world programming activities. As detailed in Algorithm  1, it operates via two steps: incremental difference calculation and difference merging, where "Overlap" denotes intersecting regions of two editing operations.
@@ 3.2.2. Instant Difference Merge
@@ -1,3 +1,3 @@ -def Hello() +def GoodBye() print("Say") - print("Hello") + print("GoodBye") 1-| def Hello() 1+| def GoodBye() 2 | print("Say") 3-| print("Hello") 3+| print("GoodBye")
Algorithm 1 Incremental Difference Detecting
A stream of edit events $E_{stream}$, where each event $e=(C_{pre},C_{post})\in E_{stream}$ encompasses the pre- and post-modification content of code snippets.
A list of merged historical edits =$\mathcal{H}=\{\Delta_{1},\Delta_{2},\dots\}$.
Initialize $\mathcal{H}\leftarrow\emptyset$ and $\Delta_{active}\leftarrow\text{null}$.
for each incoming edit event $(C_{pre},C_{post})$ in $E_{stream}$ do
  $\Delta_{new}\leftarrow\text{Diff}(C_{pre},C_{post})$
  if $\Delta_{active}$ is null then
    $\Delta_{active}\leftarrow\Delta_{new}$
  else if $\text{Overlap}(\Delta_{active},\Delta_{new})$ then
    $\Delta_{active}\leftarrow\text{Merge}(\Delta_{active},\Delta_{new})$
  else
    Add $\Delta_{active}$ to $\mathcal{H}$
    $\Delta_{active}\leftarrow\Delta_{new}$
  end if
end for
After the stream ends, add the last non-null $\Delta_{active}$ to $\mathcal{H}$.
return the list of historical trajectories $\mathcal{H}$.
@@ 3.3. Training Dataset Construction
To mitigate this, we introduce a crucial data refinement stage. We employ an LLM as a relevance filter to analyze the semantic and causal relationship between the historical edits ($H_{T}$) and the current edit ($L_{\text{gt}}$, $E_{\text{gt}}$). Based on the LLM’s assessment, each instance is processed as follows:
@@
This filtering and relabeling process transforms the preliminary training datacset into a final, optimized training set. This refined dataset not only enhances the quality of modification-oriented samples but, critically, also incorporates explicit negative samples. This dual focus is essential for training a model that is both proactive in its assistance and discerning enough to avoid generating noise.
@@ 3.4.2. Stage 2: Reinforcement Learning with DAPO
A reward of 1.0 is granted for an exact match, strongly encouraging the model to pursue perfect accuracy. If the generated edit is not an exact match but is substantially similar to the ground truth (ES > 0.5), a proportional reward is assigned. The scaling factor of 0.5 ensures this partial credit is strictly less than the reward for a perfect solution, while still guiding the model to generate semantically relevant code. A penalty of -1.0 is given for suggestions that are neither correct nor sufficiently similar, preventing the model from learning to produce irrelevant outputs.
@@ 4. Evaluation
We evaluate NES by addressing the following research questions:
@@ 4.1.2. Metrics
For RQ2, which evaluates code-editing performance, the three key metrics are: Edit Similarity (ES) for semantic alignment between the generated and target edits; Exact Match Rate (EMR) for syntactic precision of the generated edits for the modification task; and Accuracy (ACC) for correctly predict preservation.
@@ 4.1.3. Baseline Models
We compare NES to the following models.
  • Qwen (Yang et al., 2025), developed by Alibaba Cloud, is a renowned LLM, while Qwen-Coder (Yang et al., 2024) specializes in programming and software development tasks.
  • Claude (Anthropic, [n. d.]), created by Anthropic, has strong capabilities in coding and can assist with a wide range of programming tasks, including writing, debugging, and explaining code.
@@
We categorize these models into two groups based on their size. The models with relatively smaller sizes include Qwen3-4B, Qwen2.5-Coder-7B, and Seed-Coder-8B-Instruct. These models will undergo supervised fine-tuning (SFT) using the NES SFT and DAPO datasets. Zeta-7B, however, was not subjected to additional training, as it is already built on Qwen2.5-Coder-7B and incorporates Zeta’s own training data. Instead, we use Zeta-7B as a baseline.
@@
We did not select instruction-based tools such as CodeT5 (Wang et al., 2021), CoditT5 (Zhang et al., 2022b), Coeditor (Wei et al., 2024), and CoEdPilot (Liu et al., 2024a) as baselines, as these tools rely on explicit human instructions to operate. Since our evaluation environment and benchmarks do not account for such input, including these tools would lead to an unfair comparison. However, based on feedback from real-world developers, our approach provides an editing tool that is at least as effective, if not superior, in delivering practical assistance to users.

### v3

# NES: An Instruction-Free, Low-Latency Next Edit Suggestion Framework Powered by Learned Historical Editing Trajectories

@@ Abstract.
Code editing is a frequent yet cognitively demanding task in software development. Existing AI-powered tools often disrupt developer flow by requiring explicit natural language instructions and suffer from high latency, limiting real-world usability. We present NES (Next Edit Suggestion), an instruction-free, low-latency code editing framework that leverages learned historical editing trajectories to implicitly capture developers’ goals and coding habits. NES features a dual-model architecture: one model predicts the next edit location, and the other generates the precise code change—both without any user instruction. Trained on our open-sourced SFT and DAPO datasets 2, NES achieves state-of-the-art performance (75.6% location accuracy, 27.7% exact match rate) while delivering suggestions in under 250ms. Deployed at Ant Group, NES serves over 20,000 developers through a seamless Tab-key interaction, achieving effective acceptance rates of 51.55% for location predictions and 43.44% for edits—demonstrating its practical impact in real-world development workflows.
Code Edit, Reinforcement Learning (RL), Edit Location Prediction
journalyear: 2026copyright: ccconference: 34th ACM Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering; July 05–09, 2026; Montreal, QC, Canadabooktitle: 34th ACM Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering (FSE Companion ’26), July 05–09, 2026, Montreal, QC, Canadadoi: 10.1145/3803437.3805244isbn: 979-8-4007-2636-1/2026/07ccs: Software and its engineering Software development techniquesccs: Computing methodologies Machine learning
@@ 1. Introduction
With Large Language Models (LLMs) demonstrating remarkable capabilities in various aspects of software development (Di et al., 2023; Team, 2025b; Wang et al., 2024b; Guo et al., 2024; Liang et al., 2024; Rozière et al., 2023; Ding et al., 2024; Codefuse et al., 2025), AI-powered development tools such as GitHub Copilot (Copilot, 2023), Cursor (Cursor, 2024) and Zeta (Zed, 2025) are revolutionizing modern software development workflows by leveraging LLMs for code editing tasks (Wang et al., 2021; Zhang et al., 2022a; Gupta et al., 2023; Wei et al., 2024; Liu et al., 2024a). Empirical studies further highlight that code editing, which includes activities such as modifying (Li et al., 2023; Guo et al., 2025), refactoring, and maintaining existing code, is one of the most frequent and critical activities throughout the lifecycle of a software project (Nguyen et al., 2013). In particular, code completion task can be formally regarded as a restricted instance of code editing, constrained to insertions immediately following the cursor position within a partially written syntactic construct (e.g., statement, expression, or function body). The importance of efficient and accurate code editing cannot be overstated, as it directly impacts the productivity and quality of software development.
Existing research efforts have made significant advancements, particularly in translating explicit natural language instructions into code edits. However, these approaches face critical limitations that hinder their effective integration into a developer’s workflow. On one hand, most existing methods, such as CodeT5 (Wang et al., 2021), CoditT5 (Zhang et al., 2022a), Coeditor (Wei et al., 2024), CoEdPilot (Liu et al., 2024a), EfficientEdit (Wang et al., 2025) and so on, rely heavily on human instruction input to maintain a coherent reference chain, disrupting programming flow and increasing cognitive load. Additionally, tools that fail to capture users’ habitual coding behavior often generate suggestions misaligned with their natural coding style. On the other hand, code editing tasks are inherently time-sensitive, with users expecting feedback within 1s(Card, 2018). But existing tools (Team, 2025b, a; Codefuse et al., 2025; Anthropic, 2024; Guo et al., 2024) rely on computationally intensive processes (e.g., search, text understanding, reasoning), creating a tradeoff between low-latency and precise suggestions. This significant latency undermines real-time interaction and degrades user experience. Overcoming these limitations can undoubtedly enhance code editing in real-world development. However, it necessitates modifying the method of acquiring cues, which poses significant challenges in accurately identifying the editing location and determining the content within the desired timeframe.
@@
We propose NES (Next Edit Suggestion), an LLM-driven code editing framework designed for an instruction-free and low-latency (feel like zero-latency) experience. Comprising a synergistic dual-model architecture trained on the proposed Supervised Fine-Tuning (SFT) (Wei et al., 2022) and Dynamic sAmpling Policy Optimization (DAPO) (Yu et al., 2025) datasets, NES adapts to user habits, integrates seamlessly into workflows, and enhances productivity by implicitly learning developer intent.
@@
  • NES-Edit Model analyzes the developer’s editing history, including refactoring patterns and API call sequences, to generate personalized and precise code modifications for the current edit.
To improve user experience, NES leverages Prefix Caching (PC) (Zheng et al., 2024) and Speculative Decoding (SD) (Leviathan et al., 2022; Chen et al., 2023) techniques to achieve substantial inference speedups.
Proven at scale with over 20,000 developers at Ant Group, NES is an industry-ready solution that streamlines complex coding tasks. Its core innovation is a continuous workflow where developers use simple Tab key sequences (TabTabTab) to perform operations like refactoring. This unique interaction model eliminates manual steps, significantly reducing cognitive load and context-switching to create a highly efficient and immersive coding experience.
We demonstrate the practical effectiveness of NES using both open-source and industrial datasets. Our model achieves SOTA performance, with the Location Model reaching 75.6% accuracy in predicting edit placement and the Edit Model delivering a 27.7% Exact Match Rate (EMR) for generating intent-aligned code. By optimizing for Nvidia L20 GPUs, we ensure efficiency with an inference time of under 250ms. Additionally, our proposed SFT and DAPO datasets significantly improve existing Qwen2.5-Coder-7B and Seed-Coder-8B-Instruct models.
@@
  • We introduce NES  [^1] Unless otherwise specified, NES refers to the Qwen3-4B+SFT+DAPO model. See Section 5 for further details., a dual-model, LLM-powered code editing framework that leverages historical code edit patterns. Operating in an instruction-free, low-latency setting, it eliminates the need for explicit natural language input from users.
  • We construct and open-source high-quality SFT and DAPO datasets [^2] A portion of the dataset is publicly available at https://huggingface.co/datasets/codefuse-ai/CodeFuse_codeedit. The remaining data will be released incrementally., which significantly boost performance. These datasets have played a critical role in achieving dramatic improvements—by orders of magnitude—in accuracy and similarity metrics across various code editing benchmarks.
  • We have integrated and deployed NES across Nvidia L20 GPUs within Ant Group, serving 20,000+ developers. The implementation of various inference optimizations ensures that NES delivers real-time responses (under 250ms), culminating in a more fluid user experience with effective acceptance rates of 43.44% for code edits and 51.55% for location prediction.
@@ 2. Motivating Example
The process involves manual navigation, repetitive edits, and significant cognitive effort to track changes. Although LLM-driven tools can assist with locating and making edits, instruction-based tools, such as CodeT5 (Wang et al., 2021), CoditT5 (Zhang et al., 2022a), Coeditor (Wei et al., 2024), and CoEdPilot (Liu et al., 2024a), disrupt the developer’s workflow by requiring precise commands and waiting for responses, introducing additional cognitive overhead. To address these challenges, an instruction-free tool is essential, which is where NES comes in.
@@
A tedious, error-prone manual process is streamlined into an efficient ”TabTabTab ” workflow. Powered by high-quality datasets, NES predicts both the location (NES-Location) and content (NES-Edit) of edits without requiring explicit instructions. This reduces cognitive load, minimizes context switching, accelerates inference and delivers an immersive development experience.
@@ 3.1. An Overview of NES Framework
  • Next Location Suggestion: The NES-Location Model leverages real-time editing trajectories ① to predict the next likely edit location. The developer can rapidly jump to this position by pressing the Tab key or simply disregard the suggestion ②.
  • Code Edit Suggestion: Once at the location, the NES-Edit Model generates a code edit by integrating historical patterns with the current code context ③. This suggestion can be accepted with the Tab key, or the developer can continue editing manually ④, which in turn updates the edit history ⑤.
  • Dynamic loop: Each accepted suggestion triggers the next prediction, creating a continuous assistance loop ⑥. This process transforms complex refactoring tasks into a fluid Tab sequence, significantly boosting developer productivity.
@@
Incremental Difference Detection
We developed a real-time code change monitor based on incremental computation. This framework accurately captures each developer editing operation, thereby constructing a comprehensive and detailed historical editing trajectory. These trajectory data serve as crucial inputs for subsequent model training and prediction processes.
High-Quality Training Dataset Construction
Utilizing the collected authentic historical editing trajectories, we constructed a high-quality training dataset. This dataset encompasses not only sequences of editing operations but also contextual information for each edit. Therefore, this dataset can enable the training of both the NES-Location Model and the NES-Edit Model with nuanced understanding of editing patterns and their surrounding code environments.
Model Training
To endow NES with the capability for location and code editing suggestion, we designed a two-stage training methodology. Initially, both the NES-Location and NES-Edit models undergo SFT (Ouyang et al., 2022) on our large-scale historical editing datasets, enabling them to learn the fundamental patterns of code modification. Subsequently, we employ Decoupled Clip and DAPO (Yu et al., 2025), a reinforcement learning technique using high-quality preference data, to further refine the models [^3] DAPO is a reinforcement learning method that uses dynamic sampling of high-quality preferences to refine the policy, while Decoupled Clip stabilizes training by decoupling the clipping mechanism in policy updates.. This second stage aligns their behavior more closely with real-world developer intent and utility, significantly boosting the practical value of NES.
Model Inference Optimization
Acknowledging the constraints of inference resources in practical application environments and the stringent low-latency requirements inherent in real-time editing suggestion, we conducted in-depth optimization of the model inference process, and achieved real-time responsiveness to developer editing behaviors, ensuring the fluidity and practicality of the NES framework.
@@ 3.2. Incremental Difference Detection
We developed an efficient incremental difference detector built upon our AI-powered IDE, designed to accurately capture real-world code editing trajectories of developers during real-world programming activities. The algorithm operates in two primary steps: incremental difference calculation and difference merging, where ”Overlap” denotes intersecting regions of two editing operations.
@@ 3.2.2. Instant Difference Merge
@@ -1,3 +1,3 @@ -def Hello() +def GoodBye() print(”Say”) - print(”Hello”) + print(”GoodBye”) 1-— def Hello() 1+— def GoodBye() 2 — print(”Say”) 3-— print(”Hello”) 3+— print(”GoodBye”)
@@ 3.3. Training Dataset Construction
To mitigate this, we introduce a crucial data refinement stage. We employ an LLM, like Qwen, as a relevance filter to analyze the semantic and causal relationship between the historical edits ($H_{T}$) and the current edit ($L_{\text{gt}}$, $E_{\text{gt}}$). Based on the LLM’s assessment, each instance is processed as follows:
@@
This filtering and relabeling process transforms the preliminary training dataset into a final, optimized training set. This refined dataset not only enhances the quality of modification-oriented samples but, critically, also incorporates explicit negative samples. This dual focus is essential for training a model that is both proactive in its assistance and discerning enough to avoid generating noise.
@@ 3.4.2. Stage 2: Reinforcement Learning with DAPO
A reward of 1.0 is granted for an exact match, strongly encouraging the model to pursue perfect accuracy. If the generated edit is not an exact match but is substantially similar to the ground truth (ES ¿ 0.5), a proportional reward is assigned. The scaling factor of 0.5 ensures this partial credit is strictly less than the reward for a perfect solution, while still guiding the model to generate semantically relevant code. A penalty of -1.0 is given for suggestions that are neither correct nor sufficiently similar, preventing the model from learning to produce irrelevant outputs.
@@ 4. Evaluation
We evaluate NES by addressing the following questions:
@@ 4.1.2. Metrics
For RQ2, which evaluates code-editing performance, the three key metrics are: Edit Similarity (ES) (Song et al., 2024) for semantic alignment between the generated and target edits; Exact Match Rate (EMR) for syntactic precision of the generated edits for the modification task; and Accuracy (ACC) for correctly predict preservation.
@@ 4.1.3. Baseline Models
We compare NES to the models.
  • Qwen (Team, 2025b), developed by Alibaba Cloud, is a renowned LLM, while Qwen-Coder (Team, 2024) specializes in programming and software development tasks.
  • Claude (Anthropic, 2024), created by Anthropic, has strong capabilities in coding and can assist with a wide range of programming tasks, including writing, debugging, and explaining code.
@@
  • CoEdPilot (Liu et al., 2024a) jointly trains a localization LLM (to pinpoint edits) and a generation LLM. However, the reliance on explicit human instructions constrains their overall editing efficiency.
The models with relatively smaller sizes include Qwen3-4B, Qwen2.5-Coder-7B, and Seed-Coder-8B-Instruct. These models will undergo supervised fine-tuning (SFT) using the NES SFT and DAPO datasets. Zeta-7B, however, was not subjected to additional training, as it is already built on Qwen2.5-Coder-7B and incorporates Zeta’s own training data. Instead, we use Zeta-7B as a baseline.
@@
We select CoEdPilot (Liu et al., 2024a) as the baseline from the family of explicit human-instruction-based code editing tools, alongside models like CodeT5 (Wang et al., 2021), CoditT5 (Zhang et al., 2022a), Coeditor (Wei et al., 2024), EfficientEdit (Wang et al., 2025) and so on. As our evaluation environment does not natively support CoEdPilot’s required human instruction input format, we left the corresponding input field blank during testing.